OpenStudy (anonymous):

1) if an object of mass m has velocity v, then its kinetic energy K is given by k=1/2mv^2. if v is a function of time t, use the chain rule to find a formula for dK/dt. 2)When a space shuttle is launched into space, an astronaut’s body weight decreases until a state of weightlessness is achieved. The weight, W, of a 150-lb astronaut at an altitude of x kilometers above sea level is given by W=150 (6400/ 6400+x)^2 If the space shuttle is moving away from the earth’s surface at a rate of 6 km/sec, at what rate is W decreasing when x = 1000 km? (Hint: Keep in mind that x is changing over time. You can think of x as x(t), a function of time.)

OpenStudy (anonymous):

\[ {d K\over d t}=\left({1\over 2}m\right)(2v){d v\over d t} \]

OpenStudy (anonymous):

for 2, it is given that \[{d x\over d t}=-6{\rm km/s}\]the negative sign because it is moving away. apply chain rule to find the derivative of "W"

OpenStudy (anonymous):

that's what giving me problems. i have trouble applying the chain rule given that i don't quite understand it that well :/

OpenStudy (anonymous):

try it. you take the derivative of functions till you cannot \[W=150\left(6400\over6400+x\right)^2 \] here first let us make \[X={6400\over6400+x}\\ W=150X^2 \] what is the derivative of W now?

OpenStudy (anonymous):

300x??

OpenStudy (anonymous):

good. \[W=300X\] replace X and multiply with its derivative \[ W=300\left(6400\over6400+x\right)\times{d\over dx}\left(6400\over6400+x\right) \]

OpenStudy (anonymous):

x would be 1000 km , right?

OpenStudy (anonymous):

not yet. first evaluate the derivative

OpenStudy (anonymous):

would the result be 300 (x+1)?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

math cannot be solved by guess work. it is very mechanical. just follow the steps. do not think.